# A finite volume Simo-Reissner beam method for moored floating body dynamics

### v1

@@ 2.3 Rigid body motion dynamics
The motion of the rigid floating body is determined using Newton’s second law of motion, which governs the translational and rotational dynamics of the body. The external forces and moments considered include hydrodynamic forces, hydrostatic forces, gravitational forces, mooring line forces, and any additional damping or restoring forces that may be present. All these forces except gravity are integrated over the body’s surface to account for the influence of the surrounding fluid, while gravity is treated as a body force and integrated over the volume of the body. To ensure stability and accuracy, the equations of motion are solved in a coupled manner, fully accounting for the interaction between the rigid body and the surrounding fluid. This approach captures the mutual influence of fluid forces on the body’s motion and the body’s motion on the fluid flow. In OpenFOAM, the sixDoFRigidBodyMotion library is used to solve the motion of a body in six degrees of freedom (DoF). The conservation of linear and angular momentum, with rotation occurring around the centre of mass (COM), leads to the following equations for the floating body,
@@
$\frac{\partial\bm{\omega}_{rb}}{\partial t}=\mathbf{I}^{-1}_{rb}\cdot[\mathbf{M}_{rb}-\omega_{rb}\times(\mathbf{I}_{rb}\omega_{rb})]$ (20)
@@
[…]
In which $\mathbf{\mathbb{I}}$ denotes the $3\times 3$ identity matrix, $p$ is the normal pressure acting perpendicular to the surface of the floating body, and $\bm{\tau}$ represents the viscous stress term, which is the tangential force due to the fluid’s viscosity, opposing the motion of the body, $\mathbf{S}$ is rigid body surface while $\mathbf{F}_{moor}$ is the mooring restraining force, calculated by the beam solver described in Section 2.2, $r_{1}$, $r_{2}$, $r_{3}$ represent the moment arms of the hydrodynamic force, mooring restraining force, and gravitational force acting on the floating object, respectively.
@@ 2.4 Morphing mesh method
In the present study, the deforming mesh and dynamic motion of the floating rigid body is controlled using a modified version of the morphing mesh suggested by Palm and Eskilsson [53], in which the state of the floating rigid body is defined by its position, $\mathbf{B}(t)$, and orientation, $\mathbf{q}(t)$. The orientation is expressed in quaternion notation, assuming that the body has the initial orientation of $\mathbf{q}_{0}$. For an arbitrary point $\mathbf{P}_{i}$ in the CFD mesh (see Figure 3), treated as a pure quaternion with zero scalar part for the purpose of rotation, the point displacement $\bm{\delta p}_{i}$ can be computed as described in [53].
@@
where $\beta_{i}$ is the interpolation weight of point $i$, $\bm{\alpha}_{i}=[\alpha_{i}^{(x)},\alpha_{i}^{(y)},\alpha_{i}^{(z)}]^{\top}$ is the vector of scaling factors for translational motions in surge ($x$), sway ($y$), and heave ($z$), respectively, and $\mathbf{b}=[b_{x},b_{y},b_{z}]^{\top}$ represents the body displacement vector. The operator $\odot$ denotes element-wise (Hadamard) multiplication, the superscript asterisk (*) indicates the quaternion conjugate, and $W(\mathbf{P}{i})$ is the minimum distance from $\mathbf{P}{i}$ to the surface of the body. Two parameters $r_{in}$ and $r_{out}$ have been employed to define the inner and outer region, respectively. Within the $r_{in}$ region, $\mathbf{P}_{i}$ rigidly follows the body, While for the points outside the outer region ($W(\mathbf{P}_{i})>r_{out}$), the mesh remains static.
@@
The native mesh morphing class in OpenFOAM results in poor mesh quality and grid distortion due to high-amplitude rotations (see Figure LABEL:fig:omm), while the modified mesh morphing (Eqns. 23, 24 and 25) improves upon the original by adding an additional mesh morphing region to accommodate one or two translational degrees of freedom, specifically surge and/or heave. As a result, the morphing region defined by $r_{in}$ and $r_{out}$ translates the rigid body horizontally. By excluding translational motion from the inner deformation zone, the method generates deformation patterns that accurately reflect the combined effects of heave and pitch, without introducing spurious distortions (see Figure LABEL:fig:mmm). This improvement enhances the stability of numerical simulations, allowing for the successful simulation of the box’s response to the incoming wave. For a detailed discussion, see Palm and Eskilsson [53].
@@ 2.5 Coupling Strategy
Within each PIMPLE iteration, the beam equations are solved first using the position of the floating body from the previous iteration (or from the previous time step, in the first PIMPLE iteration). The beam solver calculates the mooring forces and updates the position and velocity of the mooring lines. These restraining forces and moments are applied at the fairlead location and passed to the six-degree-of-freedom (6DOF) solver to update the acceleration of the floating body. The motion of the floating body is then solved by integrating the equations of motion (Eqns. 19, 21). Based on the updated position, the computational mesh is deformed using the morphing strategy described in Eqn. 23. The position of the water-air interface is then updated by solving the advection equation for the volume fraction (Eqn. 5). Finally, the velocity-pressure coupling is resolved within the PISO sub-loop, including updates to turbulence quantities if applicable. Each time step consists of 3 PIMPLE iterations to update the fluid solver, the floating body motion, and the mooring system state. A summary of this coupling strategy is shown in Figure 5.
@@ 3 Benchmark Case
To validate the proposed approach, a numerical model is created based on the experimental work of Wu et al. [50]. This experimental setup was part of the European MaRINET2-EsflOWC project [54], in which responses and motions of mooring lines attached to a solid floating box are investigated. Details of the numerical model will be provided later in this section, once the experimental setup is described.
@@ 3.1 Experimental Setup
The experimental model consists of a box moored symmetrically to the bottom of a $30.0$ m long, $1.0$ m wide, and $1.2$ m high wave flume belonging to the Coastal Engineering Research Group at the Department of Civil Engineering of Ghent University. During this project, several tests with different wave conditions were carried out. The rigid box made of light PVC material is $0.2$ m long, $0.2$ m wide, and $0.132$ m high and has a $3.148$ kg mass. To track the box’s motion, a $0.324$ m plate was attached to the front face of the box, facing the incoming waves, on which reflective markers were installed. These reflective markers were used by a 6-DOF optical motion tracking system. Initially, the box had a draft length of $0.0786$ m, which meant 60% of its height was submerged. Figure 6 depicts a schematic view of the experimental domain.
Figure 6: Overview of the experimental domain, with both images taken from the experimental work by Wu et al. [50]: (a) top view of the domain, including the locations of wave gauges, and (b) floating box with mooring lines and attached plate.
The mooring system employed in this experiment includes four symmetrical iron chains initially slacked and postured in a catenary layout. Four load cells have been mounted on the anchor end of the mooring lines to record the tensions. On the fairlead side, each line is connected at a point 0.5 cm above the box’s corner. [22]. A detailed list of box and mooring line properties is provided in Table 1.
Table 1: Properties experimentally tested box model with mooring lines [9]
@@
Table 2: Fairleads and anchors coordinates as reported in Wu et al. [50]
@@
Fairlead 1 $(-0.1,0.1,0.0736)$
Fairlead 2 $(-0.1,-0.1,-0.0736)$
Fairlead 3 $(0.1,0.1,-0.0736)$
Fairlead 4 $(0.1,-0.1,-0.0736)$
Anchor 1 $(-1.385,0.423,-0.5)$
Anchor 2 $(-1.385,-0.423,-0.5)$
Anchor 3 $(1.385,0.423,-0.5)$
Anchor 4 $(1.385,-0.423,-0.5)$
Table 3: Wave gauges coordinates within the wave tank as reported in Wu et al. [50]
@@
Wave gauge 1 $(-2.74,0)$
Wave gauge 2 $(-0.05,0.26)$
Wave gauge 3 $(0.07,-0.36)$
@@ 3.2 Numerical model
Figure 7 gives an overview of the numerical setup and the boundary conditions used in the numerical model. Following the approach employed in Chen and Hall [9], rather than simulating the wave absorption beach in the physical wave flume, a cell stretching zone is designated in the numerical flume’s last 3 meters, which decreases the total grid count and aids in wave damping. The numerical grid is represented by a $10$ m $\times$ $1$ m $\times$ $0.957$ m (L $\times$ W $\times$ H) domain. Following the approach outlined in Domínguez et al. [22], the numerical domain used in this study is shorter than the experimental domain to enhance computational efficiency. Based on the grid convergence study reported by Chen and Hall [9], a uniform grid (expansion ratio = $1$) with the size of $0.01$, $0.01$ and $0.011$ is applied in $x$,$y$ and the $z$ direction for the 70% of the entrance length $x\in[-3.5,3.5]$ using multi-grading functionality of OpenFOAM that can divide a block in a given direction and apply different grading within each division. A graded mesh with an expansion ratio of $20$ was applied to the last 30% of the domain. The resulting fluid mesh consists of approximately 5.8 million hexahedral cells. The structure of the numerical mesh is visualised in Figure 8.
@@
Figure 7: Numerical model of the European MaRINET2 EsflOWC [54] project benchmark case
@@
Following previous numerical studies [22], a Young’s modulus of $E=2.78$ MPa was used for the mooring lines, resulting in an axial stiffness of $EA=19$ N. Each mooring line was discretised into $60$ cells. The perpendicular and tangential added mass coefficients have been set to $1.6$ and 00, respectively [22]. The drag coefficients in perpendicular and tangential directions have been set to $1.6$ and $0.5$, respectively [22]. For the normal component of ground contact with the seabed, A penalty stiffness of $1000\,\mathrm{N/m^{2}}$ and a damping coefficient of $1\,\mathrm{N\cdot s/m^{2}}$ are applied in the model, with a tangential friction coefficient of $0.01$; these values were chosen to significantly reduce penetration without compromising convergence.
In order to initialise the mooring lines to have the correct initial posture and also match the pretension observed in the experimental work [50], a straight and inclined beam starting from the relevant anchor point and pointing toward the fairlead is considered. The beam equation is solved in a separate beam simulation using the beam solver, where the anchor end of the beam is constrained with fixed displacement but free to rotate, and a displacement boundary condition is applied to the opposite end. This displacement condition is defined by a vector extending from the free end of the beam to the fairlead location. Once the beam adopts its equilibrium posture, the resulting configuration is mapped to the coupled CFD framework and used as the initial condition for the beams. The initialisation process is illustrated in Figure 9, and results in an initial pretension of $0.38$ N in each mooring line.
@@
A second-order Stokes wave generation and absorption are applied to the inlet and outlet boundaries, respectively. The parameters of the regular wave generation are listed in Table 4, following the naming convention proposed by Chen and Hall [9]. For both cases, the water depth is $0.5$ m.
@@
Case name Wave height (m) Wave period (m) Wave length (m)
H12T20 $0.12$ $2$ $4.06$
H15T18 $0.15$ $1.8$ $3.57$
The convection terms in this study were discretised using a second-order central scheme with Van Leer interpolation [55], which has been used for the momentum and the volume fraction. This combination helps reduce numerical diffusion while limiting spurious oscillations near steep gradients or interfaces. Gradient and diffusion operators were discretised using a linear second-order central difference scheme. The velocity gradient is computed using the cell-limited scheme, which constrains the gradient to ensure that when cell values are extrapolated to the faces, the resulting face values remain within the range of the surrounding cell values. Surface compression with a factor $c_{\alpha}=1$ was used in the simulations. The time derivative term is discretized using the Crank-Nicolson scheme with an off-centring coefficient of $0.9$. This coefficient, ranging between 00 and $1$, controls the balance between implicit and explicit weighting. A coefficient of $1$ results in a fully centred, second-order accurate scheme, while a coefficient of 00 reduces the scheme to the Euler-implicit method. For the beam solver, a first-order Euler time discretization scheme is employed. The overall simulation domain has been decomposed into $120$ sub-domains for parallel processing, where each sub-domain is assigned to a different CPU core. The adjustable time step based on a maximum momentum Courant number of $0.5$ was used for time integration. The interface Courant number also has been set to $0.5$. The momentum Courant number determines the main time step size, while the interface Courant number governs the sub-time-step used by the MULES algorithm to solve the volume fraction equation, ensuring accurate tracking of the phase interface without excessive movement per step. For the mesh morphing around the floating box, an outer distance of $0.45$ m and an inner distance of $0.05$ m have been considered. In addition, the modified morphing method utilizes a $0.9$ m region in the x-direction and a $0.1$ m region in the z-direction. The acceleration of the floating box was under-relaxed by a factor of $0.7$ to mitigate numerical instabilities in the rigid body motion procedure, providing a balance between solution accuracy and stability.
@@
$1$ $2004.9$- 5 800 00058000005\,800\,0005 800 000
$2$ $985.21$ $2$ 2 900 00029000002\,900\,0002 900 000
$4$ $495.78$ $4$ 1 450 00014500001\,450\,0001 450 000
$8$ $244.15$ $8.2$ 725 000725000725\,000725 000
$16$ $139.85$ $14.3$ 362 500362500362\,500362 500
$32$ $93.33$ $21.4$ 181 250181250181\,250181 250
$64$ $73.39$ $27.3$ 90 6259062590\,62590 625
$128$ $53.92$ $37.2$ 45 3124531245\,31245 312
$256$ $26.07$ $76.9$ 22 6562265622\,65622 656
$512$ $18.77$ $106.7$ 11 3281132811\,32811 328
$1024$ $16.35$ $122.5$ 5 66456645\,6645 664
Parallel speed-up $\frac{t_{1}}{t_{N}}$ refers to the improvement in execution time achieved by using multiple CPU cores ($t_{N}$) compared to a single core ($t_{1}$). It quantifies how effectively a parallelised task reduces computation time as more cores are added. Strong scaling, on the other hand, measures how the execution time of a fixed-size problem changes as the number of CPU cores increases. It helps evaluate the efficiency of a parallel algorithm when the problem size remains constant, highlighting how well the workload is distributed across increasing resources. Ideally, as the number of CPU cores increases, the execution time should scale linearly. i.e., for the same problem, using 100 CPU cores should result in an execution time that is 1% (or $0.01$) of the time required when using a single CPU core.
@@
Figure 10: Parallel performance scaliblity
Figure 11 shows the free surface interaction with incoming waves and mooring lines for the case H12T20. The iso-surface for $\alpha=0.5$ represents the free surface and is coloured by the horizontal flow velocity. Different snapshots represent the different instances of incoming waves so that an entire cycle of surge and heave motions can be observed.
@@
Table 6 provides a quantitative comparison of wave height amplitudes between experimental measurements and the present numerical model. These values are extracted using FFT analysis over the interval [6 s, 14 s], where the amplitude is calculated by identifying the dominant frequency in the spectrum and taking the corresponding magnitude, which is corrected for windowing effects and the length of the signal. The simulation shows strong agreement with the experimental amplitudes, particularly near the floating body, where the mesh resolution is finer. At wave gauges 2 and 3, the differences between the present model and the experimental results are minor. Further downstream, at wave gauges 4, 5, and 6 (Figures 13(a), 13(b), 13(c), 13(d), 13(e), and 13(f)), a gradual increase in discrepancy is evident. This can be primarily attributed to mesh coarsening in that region, which limits the model’s ability to accurately capture smaller-scale surface variations. In addition, part of the observed differences can be due to experimental uncertainties, such as wave reflections from boundaries and interactions with tank walls that are challenging to replicate in the numerical setup. Overall, the average amplitude difference compared to the experimental values is 8.39% for H12T20, with the largest deviation of 13.07% observed at wave gauge 6. For H15T18, the average difference increases slightly to 9.81%, with a maximum deviation of 13.45% also occurring at wave gauge 6. These results indicate a consistent trend of increasing discrepancy with distance from the body. Similar discrepancies have been reported in previous studies using this benchmark. For example, Chen and Hall [9] and Domínguez et al. [22] observed similar amplitude deviations in downstream gauges. This highlights common challenges in maintaining wave accuracy over distance rather than limitations specific to the present model.

### v2

@@ 2.3 Rigid body motion dynamics
The motion of the rigid floating body is determined using Newton’s second law of motion, which governs the translational and rotational dynamics of the body. The external forces and moments considered include hydrodynamic forces, hydrostatic forces, gravitational forces, mooring line forces, and any additional damping or restoring forces that may be present. All these forces except gravity and mooring forces are integrated over the body’s surface to account for the influence of the surrounding fluid, while gravity is treated as a body force and integrated over the volume of the body. Mooring loads are applied as concentrated external forces and moments at the attachment points. To ensure stability and accuracy, the equations of motion are solved in a coupled manner, fully accounting for the interaction between the rigid body and the surrounding fluid. This approach captures the mutual influence of fluid forces on the body’s motion and the body’s motion on the fluid flow. In OpenFOAM, the sixDoFRigidBodyMotion library is used to solve the motion of a body in six degrees of freedom (DoF). The conservation of linear and angular momentum, with rotation occurring around the centre of mass (COM), leads to the following equations for the floating body,
@@
$\frac{\partial\bm{\omega}_{rb}}{\partial t}=\mathbf{I}^{-1}_{rb}\cdot[\mathbf{M}_{rb}-\bm{\omega}_{rb}\times(\mathbf{I}_{rb}\bm{\omega}_{rb})]$ (20)
@@
$\mathbf{F}_{rb}=\oint_{S}(p\mathbf{\mathbb{I}}+\bm{\tau})\cdot d\mathbf{S}\;+\;\mathbf{F}_{moor}\;+\;m_{f}\mathbf{g}$ (21)
$\mathbf{M}_{rb}=\oint_{S}\mathbf{r}_{1}\times(p\mathbf{\mathbb{I}}+\bm{\tau})\cdot d\mathbf{S}\;+\;\mathbf{r}_{2}\times\mathbf{F}_{moor}\;+\;\mathbf{r}_{3}\times m_{f}\mathbf{g}$ (22)
In which $\mathbf{\mathbb{I}}$ denotes the $3\times 3$ identity matrix, $p$ is the normal pressure acting perpendicular to the surface of the floating body, and $\bm{\tau}$ represents the viscous stress term, which is the tangential force due to the fluid’s viscosity, opposing the motion of the body, $\mathbf{S}$ is rigid body surface while $\mathbf{F}_{moor}$ is the resultant restraining force transmitted by the mooring system to the floating body, obtained from the beam solver described in Section 2.2. Note that this is distinct from $\mathbf{f}_{ext}$ in the beam equations, which denotes the distributed external load acting along the mooring line, $\mathbf{r}_{1}$, $\mathbf{r}_{2}$, $\mathbf{r}_{3}$ represent the moment arms of the hydrodynamic force, mooring restraining force, and gravitational force acting on the floating object, respectively.
@@ 2.4 Morphing mesh method
In the present study, the deforming mesh and dynamic motion of the floating rigid body is controlled using a modified version of the morphing mesh suggested by palm_facilitating_2022, in which the state of the floating rigid body is defined by its position, $\mathbf{B}(t)$, and orientation, $\mathbf{q}(t)$. The orientation is expressed in quaternion notation, assuming that the body has the initial orientation of $\mathbf{q}_{0}$. For an arbitrary point $\mathbf{P}_{i}$ in the CFD mesh (see Figure 3), treated as a pure quaternion with zero scalar part for the purpose of rotation, the point displacement $\bm{\delta p}_{i}$ can be computed as described in [palm_facilitating_2022].
@@
where $\beta_{i}$ is the interpolation weight of point $i$, $\bm{\alpha}_{i}=[\alpha_{i}^{(x)},\alpha_{i}^{(y)},\alpha_{i}^{(z)}]^{\top}$ is the vector of scaling factors for translational motions in surge ($x$), sway ($y$), and heave ($z$), respectively, and $\mathbf{b}=[b_{x},b_{y},b_{z}]^{\top}$ represents the body displacement vector. The operator $\odot$ denotes element-wise (Hadamard) multiplication, the superscript asterisk (*) indicates the quaternion conjugate, and $W(\mathbf{P}{i})$ is the minimum distance from $\mathbf{P}{i}$ to the surface of the body. Two parameters $r_{in}$ and $r_{out}$ have been employed to define the inner and outer regions, respectively. Within the $r_{in}$ region, $\mathbf{P}_{i}$ rigidly follows the body, while for the points outside the outer region ($W(\mathbf{P}_{i})>r_{out}$), the mesh remains static.
@@
The native mesh morphing class in OpenFOAM results in poor mesh quality and grid distortion due to high-amplitude rotations (see Figure LABEL:fig:omm), while the modified mesh morphing (Eqns. 23, 24 and 25) improves upon the original by adding an additional mesh morphing region to accommodate one or two translational degrees of freedom, specifically surge and/or heave. As a result, the morphing region defined by $r_{in}$ and $r_{out}$ translates the rigid body horizontally. By excluding translational motion from the inner deformation zone, the method generates deformation patterns that accurately reflect the combined effects of heave and pitch, without introducing spurious distortions (see Figure LABEL:fig:mmm). This improvement enhances the stability of numerical simulations, allowing for the successful simulation of the box’s response to the incoming wave. For a detailed discussion, see palm_facilitating_2022.
@@ 2.5 Coupling Strategy
Within each PIMPLE iteration, the beam equations are solved first using the position of the floating body from the previous iteration (or from the previous time step, in the first PIMPLE iteration). The beam solver calculates the mooring forces and updates the position and velocity of the mooring lines. These restraining forces and moments are applied at the fairlead location and passed to the six-degree-of-freedom (6DOF) solver to update the acceleration of the floating body. The motion of the floating body is then solved by integrating the equations of motion (Eqns. 19, 20). Based on the updated position, the computational mesh is deformed using the morphing strategy described in Eqn. 23. The position of the water-air interface is then updated by solving the advection equation for the volume fraction (Eqn. 5). Finally, the velocity-pressure coupling is resolved within the PISO sub-loop, including updates to turbulence quantities if applicable. Each time step consists of 3 PIMPLE iterations to update the fluid solver, the floating body motion, and the mooring system state. A summary of this coupling strategy is shown in Figure 5.
@@ 3 Benchmark Case
To validate the proposed approach, a numerical model is created based on the experimental work of wu_experimental_2019. This experimental setup was part of the European MaRINET2-EsflOWC project [kisacik_efficiency_2020], in which responses and motions of mooring lines attached to a solid floating box are investigated. Details of the numerical model will be provided later in this section, once the experimental setup is described.
@@ 3.1 Experimental Setup
The experimental model consists of a box constructed from lightweight PVC, moored symmetrically to the bottom of a $30.0$ m long, $1.0$ m wide, and $1.2$ m high wave flume belonging to the Coastal Engineering Research Group at the Department of Civil Engineering of Ghent University. During this project, several tests with different wave conditions were carried out. The rigid box made of light PVC material is $0.2$ m long, $0.2$ m wide, and $0.132$ m high and has a $3.148$ kg mass. To track the box’s motion, a $0.324$ m plate was attached to the front face of the box, facing the incoming waves, on which reflective markers were installed. The mass of the plate and markers is included in the reported properties of the box (Table 1), and their influence is further discussed in Section 3.3. These reflective markers were used by a 6-DOF optical motion tracking system. Initially, the box had a draft length of $0.0786$ m, which meant 60% of its height was submerged. Figure 6 depicts a schematic view of the experimental domain.
Figure 6: Overview of the experimental domain, with both images taken from the experimental work by wu_experimental_2019: (a) top view of the domain, including the locations of wave gauges, and (b) floating box with mooring lines and attached plate.
The mooring system employed in this experiment includes four symmetrical iron chains initially slack and postured in a catenary layout. Four load cells have been mounted on the anchor end of the mooring lines to record the tensions. On the fairlead side, each line is connected at a point 0.5 cm above the box’s corner. [dominguez_sph_2019]. A detailed list of box and mooring line properties is provided in Table 1. The reported centre of gravity values in Table 1 already account for the attached wooden plate ($0.324$ m, front face) and reflective markers, ensuring that the listed values correspond to the full experimental model.
Table 1: Properties experimentally tested box model with mooring lines [chen_cfd_2022]
@@
Mooring line density $5782.10$ kg m-3
Mooring line Young’s modulus $1.82$ MPa
Mooring line Poisson’s ratio $0.5$ -
@@
Table 2: Fairleads and anchors coordinates as reported in wu_experimental_2019
@@
Fairlead 1 (0.1,0.1,0.0736)(-0.1,0.1,0.0736)
Fairlead 2 (0.1,0.1,0.0736)(-0.1,-0.1,-0.0736)
Fairlead 3 (0.1,0.1,0.0736)(0.1,0.1,-0.0736)
Fairlead 4 (0.1,0.1,0.0736)(0.1,-0.1,-0.0736)
Anchor 1 (1.385,0.423,0.5)(-1.385,0.423,-0.5)
Anchor 2 (1.385,0.423,0.5)(-1.385,-0.423,-0.5)
Anchor 3 (1.385,0.423,0.5)(1.385,0.423,-0.5)
Anchor 4 (1.385,0.423,0.5)(1.385,-0.423,-0.5)
Table 3: Wave gauges coordinates within the wave tank as reported in wu_experimental_2019
@@
Wave gauge 1 (2.74,0)(-2.74,0)
Wave gauge 2 (0.05,0.26)(-0.05,0.26)
Wave gauge 3 (0.07,0.36)(0.07,-0.36)
@@ 3.2 Numerical model
Figure 7 gives an overview of the numerical setup and the boundary conditions used in the numerical model. Following the approach employed in chen_cfd_2022, rather than simulating the wave absorption beach in the physical wave flume, a cell stretching zone is designated in the numerical flume’s last 3 meters, which decreases the total grid count and aids in wave damping. The numerical grid is represented by a $10$ m $\times$ 11 m $\times$ $0.957$ m (L $\times$ W $\times$ H) domain. Following the approach outlined in dominguez_sph_2019, the numerical domain used in this study is shorter than the experimental domain to enhance computational efficiency. A grid convergence study was performed to identify an appropriate mesh resolution that balances accuracy with computational efficiency (see Appendix A for details). Based on this study, a uniform grid (expansion ratio = $1$) with the size of $0.01$, $0.01$ and $0.011$ is applied in $x$,$y$ and the $z$ direction for the 70% of the entrance length $x\in[-3.5,3.5]$ using multi-grading functionality of OpenFOAM that can divide a block in a given direction and apply different grading within each division. A graded mesh with an expansion ratio of $20$ was applied to the last 30% of the domain. The resulting fluid mesh consists of approximately 5.8 million hexahedral cells. The structure of the numerical mesh is visualised in Figure 8.
@@
Figure 7: Numerical model of the European MaRINET2 EsflOWC [kisacik_efficiency_2020] project benchmark case
@@
Following previous numerical studies [dominguez_sph_2019], a Young’s modulus of $E=2.78$ MPa was used for the mooring lines, resulting in an axial stiffness of $EA=19$ N. Each mooring line was discretised into $60$ cells. The perpendicular and tangential added mass coefficients have been set to $1.6$ and 0, respectively [dominguez_sph_2019]. The drag coefficients in perpendicular and tangential directions have been set to $1.6$ and $0.5$, respectively [dominguez_sph_2019]. For the normal component of ground contact with the seabed, A penalty stiffness of $1000\,\mathrm{N/m^{2}}$ and a damping coefficient of $1\,\mathrm{N\cdot s/m^{2}}$ are applied in the model, with a tangential friction coefficient of $0.01$; these values were chosen to significantly reduce penetration without compromising convergence.
In order to initialise the mooring lines to have the correct initial posture and also match the pretension observed in the experimental work [wu_experimental_2019], a straight and inclined beam starting from the relevant anchor point and pointing toward the fairlead is considered. The beam equation is solved in a separate beam simulation using the beam solver, where the anchor end of the beam is constrained with fixed displacement but free to rotate, and a displacement boundary condition is applied to the opposite end. This displacement condition is defined by a vector extending from the free end of the beam to the fairlead location. Once the beam adopts its equilibrium posture, the resulting configuration is mapped to the coupled CFD framework and used as the initial condition for the beams. The initialisation process is illustrated in Figure 9, and results in an initial pretension of $0.38$ N in each mooring line.
@@
A second-order Stokes wave generation and absorption are applied to the inlet and outlet boundaries, respectively. The parameters of the regular wave generation are listed in Table 4, following the naming convention proposed by Chen and Hall [chen_cfd_2022]. For both cases, the water depth is $0.5$ m.
@@
Case name Nominal wave height (m) FFT‐based wave height (m) Wave period (s) Wavelength (m)
H12T20 0.12 0.1088 2.0 4.06
H15T18 0.15 0.1208 1.8 3.57
Rather than using the nominal heights 0.12 m (H12T20) and 0.15 m (H15T18) straight from the test report, the experimental wave-gauge signals were first analysed with a fast Fourier transform. The magnitude of the dominant spectral peak was taken as the true regular-wave height. This gives $0.1088$ m for H12T20 and $0.1208$ m for H15T18. These FFT-based heights were then prescribed at the inlet so that the simulated waves match the measured free-surface records.
The convection terms in this study were discretised using a second-order central scheme with Van Leer interpolation [van_leer_towards_1974], which has been used for the momentum and the volume fraction. This combination helps reduce numerical diffusion while limiting spurious oscillations near steep gradients or interfaces. Gradient and diffusion operators were discretised using a linear second-order central difference scheme. The velocity gradient is computed using the cell-limited scheme, which constrains the gradient to ensure that when cell values are extrapolated to the faces, the resulting face values remain within the range of the surrounding cell values. The simulations use a surface-compression coefficient of $c_{\alpha}=1$. The time-derivative term is discretised using the Crank-Nicolson scheme with an off-centring coefficient of $0.9$. This coefficient, ranging between 0 and $1$, controls the balance between implicit and explicit weighting. A coefficient of $1$ results in a fully centred, second-order accurate scheme, while a coefficient of 0 reduces the scheme to the Euler-implicit method. For the beam solver, a first-order Euler time discretisation scheme is employed. The overall simulation domain has been decomposed into $120$ sub-domains for parallel processing, where each sub-domain is assigned to a different CPU core. The adjustable time step based on a maximum momentum Courant number of $0.5$ was used for time integration, as detailed in Appendix B. The interface Courant number has also been set to $0.5$. The momentum Courant number determines the main time step size, while the interface Courant number governs the sub-time-step used by the MULES algorithm to solve the volume fraction equation, ensuring accurate tracking of the phase interface without excessive movement per step. For the mesh morphing around the floating box, an outer distance of $0.45$ m and an inner distance of $0.05$ m have been considered. In addition, the modified morphing method utilises a $0.9$ m region in the x-direction and a $0.1$ m region in the z-direction. The acceleration of the floating box was under-relaxed by a factor of $0.7$ to mitigate numerical instabilities in the rigid body motion procedure, providing a balance between solution accuracy and stability.
@@
11 $2004.9$ - 5 800 0005\,800\,000
22 $985.21$ 22 2 900 0002\,900\,000
44 $495.78$ 44 1 450 0001\,450\,000
88 $244.15$ $8.2$ 725 000725\,000
$16$ $139.85$ $14.3$ 362 500362\,500
$32$ $93.33$ $21.4$ 181 250181\,250
$64$ $73.39$ $27.3$ 90 62590\,625
$128$ $53.92$ $37.2$ 45 31245\,312
$256$ $26.07$ $76.9$ 22 65622\,656
$512$ $18.77$ $106.7$ 11 32811\,328
$1024$ $16.35$ $122.5$ 5 6645\,664
Parallel speed-up $\frac{t_{1}}{t_{N}}$ refers to the improvement in execution time achieved by using multiple CPU cores ($t_{N}$) compared to a single core ($t_{1}$). It quantifies how effectively a parallelised task reduces computation time as more cores are added. Strong scaling, on the other hand, measures how the execution time of a fixed-size problem changes as the number of CPU cores increases. It helps evaluate the efficiency of a parallel algorithm when the problem size remains constant, highlighting how well the workload is distributed across increasing resources. Ideally, as the number of CPU cores increases, the execution time should scale linearly; i.e., for the same problem, using 100 CPU cores should result in an execution time that is 1% (or $0.01$) of the time required when using a single CPU core.
@@
Figure 10: Parallel performance scalability
Figure 11 shows the free surface interaction with incoming waves and mooring lines for the case H12T20. The iso-surface for $\alpha=0.5$ represents the free surface and is coloured by the horizontal flow velocity. Each snapshot represents a successive instance of incoming waves so that an entire cycle of surge and heave motions can be observed.
@@
Table 6 provides a quantitative comparison of wave height amplitudes between experimental measurements and the present numerical model. These values are extracted using FFT analysis over the interval [6 s, 14 s], where the amplitude is calculated by identifying the dominant frequency in the spectrum and taking the corresponding magnitude, which is corrected for windowing effects and the length of the signal. The simulation shows strong agreement with the experimental amplitudes, particularly near the floating body, where the mesh resolution is finer. At wave gauges 2 and 3, the differences between the present model and the experimental results are minor. Further downstream, however, a gradual divergence appears at gauges 4–6 (Figures 13(a), 13(b), 13(c), 13(d), 13(e), and 13(f)). This trend reflects the combined influence of weak residual currents produced by partial reflections at the glass walls, higher-order harmonics introduced by the hinged-flap wavemaker that are absent from the Stokes-II field imposed in the numerical relaxation zone, and mesh coarsening in the absorption region, which increases numerical dispersion and damps the higher-frequency components of the wave train. Among these possible sources, the dominant factor is the gradual divergence at the downstream gauges, which is mainly caused by residual reflections in the experimental wave basin and numerical dispersion introduced by mesh coarsening in the absorption zone. These effects outweigh the smaller contributions from wavemaker startup transients and higher-order harmonics. Overall, the average amplitude difference compared to the experimental values is 0.86% for H12T20, with the largest deviation of 2.27% observed at wave gauge 6. For H15T18, the average difference increases slightly to 1.30%, with a maximum deviation of 4.0% also occurring at wave gauge 6. Similar discrepancies have been reported in previous studies using this benchmark. For example, chen_cfd_2022 and dominguez_sph_2019 observed similar amplitude deviations in downstream gauges. This highlights common challenges in maintaining wave accuracy over distance rather than limitations specific to the present model.